I would like to dilute 3 litres of 30% salt solution to 20% thank you. Jan
Answer by lwsshak3(11628)

I would like to dilute 3 litres of 30% salt solution to 20% thank you. Jan
***
let x=amt of water to add to 3 litres of 30% salt solution
30%*3+0%x=20%(3+x)
0.9+0=0.6+0.2x
0.2x=0.3
x=1.5
amt of water to add=1.5 litres